OpenStudy (cherry17ann):

4-Step Equation w+x+y+z=6 3w-x+y-z=-3 2w+2x-2y+z=4 2w-x-y+z=-4 **answers will be in fractions

OpenStudy (mathmale):

This is not a "four-step equation." This is a system of four equations in four unknowns that you are to "solve simultaneously." How have you solved similar problems in the past?

OpenStudy (cherry17ann):

Only three step, i kind of know how to begin but i get lost..

OpenStudy (mathmate):

And the method? Simple elimination, substitution, Gauss Elimination, Cramer's Rule? Most of the above methods apply as well to system of 3 variables as to 4, but it is likely to need many more steps with 4 variables. You'd need to work on this systematically, showing work on all steps. Yes, all of the values are fractions.

OpenStudy (dumbcow):

The most straightforward approach is to reduce system down to 3 equations, then reduce it down to 2 equations, finally to 1 equation. Then once you know 1 variable, use back-substitution to find all other variables. To reduce the system, you must eliminate a variable by combining equations. Eliminate "z" by adding equations (1) and (2). (It also gets rid of x) \[(5) 4w +2y = 3\] Eliminate "z" by subtracting (4) from (3). (It also gets rid of w) \[(6) 3x - y = 8\] Eliminate "z" by adding (2) and (4). \[(7) 5w -2x = -7\] Now we have system of 3 equations, 3 variables Continue by eliminating "y". Add (5) with 2*(6) \[(8) 4w + 6x = 19\] Now we have system of 2 variables. Continue by eliminating "x" Add 3*(7) with (8) \[19w = -2\] \[w = -\frac{2}{19}\] Now we can use back-substitution to find rest of variables. plug w into equation (8), solve for x \[4(-\frac{2}{19}) +6x = 19\] \[x = \frac{19^2 +8}{6*19} = \frac{123}{38}\] plug x into (6), solve for y: \[3(\frac{123}{38}) -y = 8\] \[y = \frac{65}{38}\] Finally, plug w,x,y into (1) . solve for z \[-\frac{2}{19} + \frac{123}{38} + \frac{65}{38} + z = 6\] \[z = \frac{228+4-123-65}{38} = \frac{22}{19}\]